Mike is at an amusement park. The admission cost is $25. Mike also wants to ride some of
the rides, which are $2 each. Write an expression that can be used to find the cost for Mike to ride r
rides.
If Mike spent $57 at the amusement park (buying only rides & admission), how many rides did she
get to go on

Answer by mananth(16946)

The admission cost is $25
rides are $2 each
Number of rides = r
cost for r rides = 2r
Total expense = 25+2r
$57 at the amusement park (buying only rides & admission)
57= 25+2r
57-25= 2r
32 =2r
r=16 number of rides